# LLMs Meet Multimodal Generation and Editing: A Survey

### v1

@@ Abstract
With the recent advancement in large language models (LLMs), there is a growing interest in combining LLMs with multimodal learning. Previous surveys of multimodal large language models (MLLMs) mainly focus on understanding. This survey elaborates on multimodal generation across different domains, including image, video, 3D, and audio, where we highlight the notable advancements with milestone works in these fields. Specifically, we exhaustively investigate the key technical components behind methods and multimodal datasets utilized in these studies. Moreover, we dig into tool-augmented multimodal agents that can use existing generative models for human-computer interaction. Lastly, we also comprehensively discuss the advancement in AI safety and investigate emerging applications as well as future prospects. Our work provides a systematic and insightful overview of multimodal generation, which is expected to advance the development of Artificial Intelligence for Generative Content (AIGC) and world models. A curated list of all related papers can be found at https://github.com/YingqingHe/Awesome-LLMs-meet-Multimodal-Generation.
@@
Multimodal Generation, Large Language Models, Text-to-Image, Text-to-Video, Text-to-3D, Text-to-Audio, Multimodal Agents, AI Safety, Diffusion Models, Transformers, Generative AI.
@@ 1 Introduction
The interactions between human beings and the physical world involve information from multiple modalities, such as language, vision, and audio. Therefore, realizing a world simulator also requires the model to be able to perceive and respond with multimodal information in a flexible manner. Recently, OpenAI proposed a foundation video generation model Sora [1] that is capable of generating highly realistic videos as world simulators. It makes great progress in simulating or generating real-world videos but is unable to generate other modalities, such as text, 3D, and audio. Also, it lacks the ability to perceive other modalities such as image, video, 3D, and audio, making it an incomprehension world simulator.
@@
In the past few years, researchers focused on the generation of each single modality and have achieved great progress: In terms of text generation, we have witnessed a qualitative leap in the effectiveness of text generation: From BERT [2], GPT1 [3], GPT2 [4], GPT3 [5], GPT4 [6] to ChatGPT [7], LLaMA [8, 9], the number of model parameters and training samples has grown rapidly, resulting in the continual growth of modal abilities and product deployment. In the visual generation field, with the rapid progress of diffusion models and large-scale image-text datasets, image generation has achieved remarkable achievement and can synthesize high-quality images based on various user-provided text prompts [10, 11, 12, 13]. Subsequently, significant advancements have been made in the field of video generation through the utilization of video diffusion models and large-scale video-language datasets. Notably, several groundbreaking works have emerged, such as [14, 15, 16, 17, 18, 19, 20, 21, 22] and Sora [1]. For the 3D generation, with the emergence of the CLIP [23] model, some methods [24, 25, 26] try to bring the text information to the rendered image of 3D representation (i.e., point cloud, mesh, NeRF [27] and gaussian splatting [28]), these approaches have led to significant developments in the text-to-3D generation. Additionally, the integration of Stable Diffusion (SD) [10] with text-to-image rendering has enabled a series of works in text-to-3D generation [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. The powerful text-to-image model helps the 3D generation achieve higher performance and better results. In the area of audio generation, a series of representative works tackle the different audio domains such as [44, 45, 46] for text-to-audio, [47, 48, 49] for text-to-music, and [50, 51, 52, 53, 54, 55] for text-to-speech, and they have achieved significant performance in generating high-quality natural sounds, music, and human-level speech.
@@
With the notable progress and significant performance improvements of Large Language Models (LLMs), other non-text modalities have started harnessing the power of LLMs to either enhance their generation pipelines or integrate text generation and non-text generation in a unified system, aiming to achieve more advanced functionalities and improved generation performance. In the context of image generation, there are two categories of approaches that have achieved significant integration with language models. The first category involves encoding visual information into discrete token IDs, trying to unify visual understanding and generation [56, 57, 58, 59, 60, 61]. Specifically, visual information is encoded into token representations, and LLMs directly comprehend and generate visual tokens, enabling simultaneous visual understanding and generation. The second category focuses on leveraging LLMs to enhance the generation quality of existing pretrained text-to-image (T2I) models: One category of works involves utilizing an LLM as a layout planner to incorporate knowledge of object spatial positions, quantity, and object size, enabling the generation of required bounding boxes [62, 63, 64, 65, 66]. After obtaining the bounding boxes, the images can be generated through a grounded text-to-image (T2I) model [67]. Another approach involves utilizing LLMs to expand input user prompts [68]: By providing highly detailed and comprehensive user prompts, LLMs generate high-quality images by enriching the prompt information. With the assistance of LLMs, visual generation has achieved higher generation quality, improved prompt following capabilities, dialogic function, and user-friendly interface. In video generation, LLMs serve as the general backbone for unified multimodal joint generation [69, 70], video layout planning [71, 72, 63, 73, 74] and temporal prompt generation [75, 76, 77, 78, 79] for dynamics guidance. For 3D generation and editing, LLMs serve as a bridge between users and 3D assets, which improves interaction efficiency [80, 81] and helps users understand [82, 83] the 3D assets. In the context of audio generation and editing, the role of language models primarily lies in serving as coordinated backbones for multimodal audio [84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96], conditioners for specific tasks [97, 98, 99], labelers for audio understanding [100, 101, 102], and agents for interactive generation/editing [103, 104, 105, 106, 107, 108], and as inspiration for novel approaches [53, 47, 109, 48, 110, 111]. The growing utilization of LLMs in the audio domain is not only transforming our way of engaging with sound and music but also expanding the boundaries at the crossroads of AGI and audio technologies. In addition, multimodal agents integrate multiple modalities into a single system to develop a universal system, that possesses the ability of comprehension and generation of non-text modalities. Therefore, LLMs are increasingly playing an indispensable role in generating various modes of content.
@@
To empower the world simulator and promote the development of multimodal generation, in this work, we provide a comprehensive review of works involving LLMs in the generation of multiple modalities and the roles of LLMs in this process. As shown in Fig. 1, we summarize the roles of LLMs into several key aspects, such as evaluator, labeller, instruction processor, planner, provider of semantic guidance, or as backbone architectures. Also, we discuss the safety issues in the AIGC era in Sec. 9, with the emerged applications, and the potential future prospects in Sec. 10 and Sec. 11.
@@
  • We present the first systematic review of LLMs applied to the generation of multiple modalities, including images, videos, 3D, and audio.
@@
  • We summarize the various roles of LLMs in the generation process in each modality from a technical view.
@@
1.1 Content Overview
We first review related surveys on modality-specific generation and LLMs in Sec. 2. We then briefly review the basic techniques of representative generative models, multi-modality encoder, transformer, and LLMs in Sec. 3. Next, we review the LLM-based visual generation in different visual modalities, including image in Sec. 4, video in Sec. 5, 3D in Sec. 6, audio in Sec. 7, and multimodal agents in Sec. 8, respectively. Lastly, we discuss the safety aspect of generative AI in Sec. 9 and several potential future directions for the LLMs-based multimodal generation field in Sec. 11.
@@ 1.2 Scope
This survey explores the generation of multiple modalities, including images, videos, 3D models, and audio. The multimodal generation in our survey includes the separate generation of different modalities as well as the joint generation of multiple modalities. We will not dig into pure text generation extensively, as there have already been many surveys specifically focusing on the advancements in that field [112, 113, 114]. Our primary focus is on how the recent emergence of large language models in the past few years can assist in the generation of other visual and audio modalities, especially in the open-domain generation. This will aid us in designing better unified generative models for multi-modalities. In detail, we focus on the following tasks:
  • Image generation and editing: Image generation aims to create various open-domain image contents, including pictures, photos, or stylized drawings from user-provided textual descriptions. Image editing aims to modify the input image content and can be based on user instructions.
  • Video generation and editing, where models generate or modify arbitrary and various dynamic visual contents guided by free-form textual descriptions;
  • 3D generation and editing, which a task for generating and editing 3D objects, scenes or avatars with user-provided textual descriptions.
  • Audio generation and editing, where textual descriptions are used to generate audio, including general sounds, music, and speech. Audio editing tasks such as adding, removing, or inpainting involve modifying existing audio using textual descriptions.
@@ 2 Related Surveys
Survey on modality-specific generation. A series of surveys focus on single modality generation, such as [115] for image generation, [116] for video generation, [117] for 3D generation, [118] for audio generation. However, the previous generation paradigm mainly employs pretrained Clip [23], Clip-related variant [119], or language encoder T5 [120] to achieve open-domain text-guided generation. With the emergence of LLMs, there are more and more generation works that leverage the powerful LLM to enhance the generation of each modality. Our work mainly discusses the role of LLMs in the generation of each modality, which is absent in previous surveys.
Survey on LLMs Numerous surveys have been conducted to explore various aspects of LLMs. For example, [121] offers a comprehensive examination of LLMs-based autonomous agents. Additionally, [122], and [123] look into multimodal LLMs (MLLMs), introducing papers that combine LLMs with other non-text modalities. They review papers for both multimodal understanding and generation in a mixed manner, introduce primarily multimodal understanding works, and focus less on multimodal generation. In contrast, our work primarily concentrates on the generation aspect, aiming to investigate the performance and functionality improvement that LLMs bring to each modality’s generation process, leading to a better AI-generated world with multiple modalities.
@@ 3 Preliminaries
In this section, we first review different types of generative models in Sec. 3.1. Then, we present multimodal alignment models in Sec. 3.2. Lastly, we summarize the foundational architecture behind large language models in Sec. 3.4.
@@ 3.1.4 Diffusion Model
The diffusion model is proposed in [127], which first gives the prototype of recent diffusion models. However, the foundational structure of the modern diffusion model, which led to a revolution of generation paradigms, was proposed in Denoised Diffusion Probabilistic Models[128]. It is elegant in training and mythical improvement and only introduced a simple regression loss. As shown in 2, the diffusion model turned the complex work into a series of denoising tasks, primarily consisting of two steps: the injection of prior noise into the data and the denoising prediction.
@@ 4.1.1 Text-guided Image Generation with CLIP
It is worth mentioning that, the application of LLMs for the image domain has been extensively investigated, especially for image understanding. LLMs can effectively serve as unified processors of visual tokens and language tokens [203, 204, 205, 206, 207, 142, 208, 209, 9, 210, 211, 212, 134, 213, 8, 214, 215, 216], tool coordinators [217, 103, 218, 219], or analysts of upstream visual model outputs [220, 221, 222]. Inspired by these works, numerous works further take advantage of LLMs for image generation, and the milestone works are shown in Figure 3. In the following, we will introduce the progress achieved in the task of image generation after the emergence of LLMs.
@@ 4.2.2 Image Editing with LLMs
MGIE [244] investigates the MLLMs to do the image editing task. The proposed MGIE can learn to convert expressive human instructions to editing guidance. The editing model is also trained to follow the editing guidance in an end-to-end way. The effectiveness of the MGIE is verified on photoshop-style manipulations, global photo optimization and local editing.
@@ 5.1.1 Text to Video Generation with CLIP
Based on the type of generative models, there are two main paradigms: one is based on diffusion models, and the other is based on transformer models combined with discrete codebooks trained by predicting the next token. Diffusion models have become a mainstream paradigm due to their ease of training. Within the diffusion framework, there are pixel-level video diffusion models [272, 16, 17] and latent-level video diffusion models [14, 15, 273, 274]. Pixel-level approaches exhibit better text alignment but require substantial computational resources. On the other hand, latent-level models are more efficient as they reduce redundancy in video data. Different works in video generation have different focuses. Some emphasize photorealistic or high-definition outputs, aiming to improve the quality of generated videos. Others focus on controllable generation, such as image-to-video approaches that enable local control over motion regions, trajectory control for object and camera movement directions, and the use of sketch, depth, and pose to control structure. Some works concentrate on generating longer videos or exploring better network architectures.
@@ 5.1.2 Text to Video Generation with LLMs
Recently, several works have also taken advantage of multimodal LLMs [69, 70] for the task of video generation. For example, VideoPoet [69] leverages a pre-trained auto-regressive transformer model to handle multimodal data for synthesizing videos with temporal consistency and high-motion fidelity. It adapts LLMs training techniques, allowing for task-specific video generation, including tasks such as text-to-video and image-to-video conversion. Another work MAGVIT-v2 [70] explores the tokenization technique of Multimodal LLMs. It transforms visual inputs into discrete tokens, enhancing Large Language Models’ performance on image and video generation tasks. It outperforms diffusion models on benchmarks like ImageNet and Kinetics, provides video compression on par with advanced codecs, and improves action recognition through effective representation learning.

### v2

@@ Abstract
With the recent advancement in large language models (LLMs), there is a growing interest in combining LLMs with multimodal learning. Previous surveys of multimodal large language models (MLLMs) mainly focus on multimodal understanding. This survey elaborates on multimodal generation and editing across various domains, comprising image, video, 3D, and audio. Specifically, we summarize the notable advancements with milestone works in these fields and categorize these studies into LLM-based and CLIP/T5-based methods. Then, we summarize the various roles of LLMs in multimodal generation and exhaustively investigate the critical technical components behind these methods and the multimodal datasets utilized in these studies. Additionally, we dig into tool-augmented multimodal agents that can leverage existing generative models for human-computer interaction. Lastly, we discuss the advancements in the generative AI safety field, investigate emerging applications, and discuss future prospects. Our work provides a systematic and insightful overview of multimodal generation and processing, which is expected to advance the development of Artificial Intelligence for Generative Content (AIGC) and world models. A curated list of all related papers can be found at https://github.com/YingqingHe/Awesome-LLMs-meet-Multimodal-Generation.
@@
LLMs, MLLMs, Multimodal Generation, Text-to-Image, Text-to-Video, Text-to-3D, Text-to-Audio, Multimodal Agents, AI Safety, Diffusion Models, Transformers, Generative AI, AIGC.
@@ 1 Introduction
The interactions between human beings and the physical world involve information from multiple modalities, such as language, vision, and audio. Therefore, realizing a world simulator also requires the model to be able to perceive and respond with multimodal information in a flexible manner. Recently, OpenAI introduced a foundation text-to-video generation model termed Sora [1] that is capable of generating highly realistic videos as world simulators. It makes great progress in simulating or generating real-world scenes but is unable to generate other modalities, such as text, 3D, and audio. Also, it lacks the ability to perceive other modalities such as image, video, 3D, and audio, making it an incomprehension world simulator.
@@
In the past few years, researchers focused on the generation of each single modality and have achieved great progress: In terms of text generation, we have witnessed a qualitative leap in the performance of natural language processing tasks: From BERT [2], GPT1 [3], GPT2 [4], GPT3 [5], GPT4 [6] to ChatGPT [7], LLaMA [8, 9], the number of model parameters and training samples has grown rapidly, resulting in the continual growth of modal abilities and product deployment. In the visual generation field, with the rapid progress of diffusion models and large-scale image-text datasets, text-to-image (T2I) generation has achieved remarkable achievement and can synthesize high-quality images based on various user-provided text prompts, such as SDXL [10] and PIXART-$\alpha$ [11]. Subsequently, significant advancements have been made in the field of text-to-video generation through the utilization of video diffusion models [12] and large-scale video-language datasets [13]. Notably, several milestone works have emerged, such as [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] and Sora [1]. For the 3D generation, with the emergence of the CLIP [24] model, some methods [25, 26, 27] try to align the text information to the rendered images from 3D representations, i.e., mesh, point cloud, NeRF [28] and gaussian splatting [29]). These approaches have led to significant developments in text-to-3D generation. Additionally, the integration of Stable Diffusion (SD) [30] with text-to-image rendering has enabled a series of works in text-to-3D generation [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. The powerful text-to-image model helps the 3D generation achieve higher performance and better visual results. In the area of text-to-audio generation, a series of representative works tackle different audio domains such as [46, 47, 48] for text-to-audio, [49, 50, 51] for text-to-music, and [52, 53, 54, 55, 56, 57] for text-to-speech, and they have achieved significant performance in generating high-quality natural sounds, music, and human-level speech.
@@
With the notable progress and significant performance improvements of Large Language Models (LLMs), other non-text modalities have started harnessing the power of LLMs to either enhance their generation quality or integrate multiple modalities in a unified system to achieve more powerful functionalities. In the context of image generation, the integration of LLMs can be divided into two categories. The first category involves encoding visual information into discrete token indices, trying to unify visual understanding and generation [58, 59, 60, 61, 62, 63]. Specifically, visual information is encoded into token representations, and LLMs directly comprehend and generate visual tokens, enabling simultaneous visual understanding and generation. The second category focuses on leveraging LLMs to enhance the generation quality of existing pretrained T2I models: One type of approach utilizes an LLM as a layout planner to incorporate knowledge of object spatial positions, quantity, and object size, enabling the generation of required bounding boxes [64, 65, 66, 67, 68]. After obtaining the bounding boxes, the images can be generated through a grounded T2I model such as GLIGEN [69]. Another approach utilizes LLMs to expand input user prompts [70]: By providing highly detailed and comprehensive prompts, LLMs generate images with high quality and richness. With the assistance of LLMs, image generation has achieved higher generation quality, improved prompt following capabilities, dialogic function, and user-friendly interface.
Similar to image domain, in video generation, LLMs serve as the general backbone for unified multimodal joint generation [71, 72], video layout planning [73, 74, 65, 75, 76] and temporal prompt generation [77, 78, 79, 80, 81] for temporal dynamics guidance. For 3D generation and editing, LLMs serve as a bridge between users and 3D assets, which improves interaction efficiency [82, 83] and helps users understand the 3D assets [84, 85]. In the context of audio generation and editing, the role of LLMs primarily lies in serving as coordinated backbones for multimodal audio [86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98], conditioners for specific tasks [99, 100, 101], labelers for audio understanding [102, 103, 104], agents for interactive generation and editing [105, 106, 107, 108, 109, 110], as well as inspiration for novel approaches [55, 49, 111, 50, 112, 113]. The growing utilization of LLMs in the audio domain is not only transforming our way of engaging with sound and music but also expanding the boundaries at the crossroads of AGI and audio technologies. Besides, multimodal agents [105, 114, 115, 116, 117, 118] integrate lots of AIGC tools into the framework as a universal system, which relies on LLMs to invoke tools but endows LLMs with the ability to comprehend and generate content of non-text modalities. Generally, LLMs significantly play an indispensable role in generating various modes of content.
@@
To promote the development of multimodal generation and empower the world simulator, in this work, we provide a comprehensive review of works involving LLMs in the generation of multiple modalities. As shown in Fig. 1, we summarize the roles of LLMs into several key aspects, such as evaluator, labeller, instruction processor, planner, provider of semantic guidance, or as backbone architectures. Additionally, we discuss the development in the generative AI safety topic in Sec. 9, with the emerged applications, and the potential future prospects in Sec. 10 and Sec. 11.
@@
  • We present the first systematic review of LLMs applied to the generation and editing of multiple modalities, including images, videos, 3D, and audio.
@@
  • We summarize the various roles of LLMs in the generation or editing process for each modality from a technical view.
@@ 1.1 Scope
This survey explores the generation of multiple modalities, including images, videos, 3D models, and audio. The multimodal generation in our survey includes the separate generation of different modalities as well as the joint generation of multiple modalities. We will not dig into pure text generation and processing extensively, as there have already been many surveys specifically focusing on the advancements in that field [119, 120, 121]. Our primary focus is on how the recent emergence of LLMs in the past few years can assist in the generation of other vision and audio modalities, especially in the open-domain generation. This will aid us in designing better unified generative models for multi-modalities. Note that the tasks and works we discussed are primarily language-based generation and editing. Unconditional generation and other non-text-based editing are not our primary focus since they are either limited to a small domain or lack flexibility and controllability. In detail, we focus on the following tasks:
  • Text-to-image generation and editing: Image generation aims to create various open-domain image contents, including pictures, photos, or stylized drawings from user-provided textual descriptions. Image editing aims to modify the input image content and can be based on user instructions.
  • Text-to-video generation and editing, where models generate or modify arbitrary and various dynamic visual contents guided by free-form text descriptions.
  • Text-to-3D generation and editing, which is a task for generating and editing 3D objects, scenes, or avatars with user-provided textual descriptions.
  • Text-to-audio generation and editing, where textual descriptions are used to generate audio, including general sounds, music, and speech. Audio editing tasks, such as adding, removing, or inpainting, can all be performed by modifying existing audio content through textual descriptions.
@@
1.2 Content Overview
We first review related surveys on both single modality generation and LLMs in Sec. 2. We then briefly review the basic techniques of representative generative models, multimodal alignment models, LLMs, and MLLMs in Sec. 3. Next, we review the LLM-based vision generation in different visual modalities, including image in Sec. 4, video in Sec. 5, 3D in Sec. 6, audio modality in Sec. 7, and multimodal agents in Sec. 8, respectively. Lastly, we review the safety aspect of generative AI in Sec. 9, emerging applications in Sec. 10, and potential future directions for the LLMs-based multimodal generation field in Sec. 11.
@@ 2 Related Surveys
Survey on modality-specific generation. A series of surveys focus on single modality generation, such as [122] for image generation, [123] for video generation, [124] for 3D generation, [125] for audio generation. However, the previous generation paradigm mainly employs pretrained CLIP [24], CLIP-related variant [126], or language encoder T5 [127] to achieve open-domain text-guided generation. With the emergence of LLMs, there is a growing trend of leveraging powerful LLMs to enhance the generation of content for each modality. Our work aims to provide a comprehensive survey on the role of LLMs in the generation of various types of modalities, an aspect that is absent from previous surveys.
Survey on LLMs and MLLMs. Numerous surveys have been conducted to explore various aspects of LLMs. For example, [128] offers a comprehensive examination of LLMs-based autonomous agents. Additionally, [129], and [130] look into MLLMs, introducing papers that combine LLMs with other non-text modalities. They review papers on multimodal understanding and generation in a mixed manner, introducing primarily multimodal understanding works and focusing less on multimodal generation. In contrast, our work primarily concentrates on the generation aspect, aiming to investigate the performance and functionality improvement that LLMs bring to each modality’s generation process, leading to a better AI-generated world with various modalities.
@@ 3 Preliminaries
In this section, we first review different types of generative models in Sec. 3.1. Then, we illustrate multimodal alignment models in Sec. 3.2. Lastly, We introduce the technical principle of large language models in Sec. 3.3, and explain multimodal large language models in Sec. 3.4
@@ 3.1.4 Diffusion Model
The diffusion model is proposed in [134], which first gives the prototype of recent diffusion models. However, the foundational structure of the modern diffusion model, which led to a revolution of generation paradigms, was proposed in Denoised Diffusion Probabilistic Models[135]. It is elegant in training and mythical improvement and only introduced a simple regression loss. As shown in Fig. 2, the diffusion model turned the complex work into a series of denoising tasks, primarily consisting of two steps: the injection of prior noise into the data and the denoising prediction.
@@ 4.1.1 Text-guided Image Generation with CLIP
It is worth mentioning that, the application of LLMs for the image domain has been extensively investigated, especially for image understanding. LLMs can effectively serve as unified processors of visual tokens and language tokens [212, 213, 214, 215, 216, 150, 217, 218, 9, 219, 220, 221, 142, 222, 8, 223, 224, 225], tool coordinators [116, 105, 117, 226], or analysts of upstream visual model outputs [227, 228, 229]. Inspired by these works, numerous works further take advantage of LLMs for image generation, and the milestone works are shown in Fig. 3. In the following, we will introduce the progress achieved in the task of image generation after the emergence of LLMs.
@@ 4.2.2 Image Editing with LLMs
MGIE [251] investigates the MLLMs to do the image editing task. The proposed MGIE can learn to convert expressive human instructions to editing guidance. The editing model is also trained to follow the editing guidance in an end-to-end way. The effectiveness of the MGIE is verified on photoshop-style manipulations, global photo optimization and local editing. SmartEdit [253] is another recent work that utilizes MLLMs for complex instruction-based image editing. SmartEdit analyzes the performance of instruction-based image editing models under complex instructions and proposes a Bidirectional Interaction Module to make the image feature output by pre-trained image encoder and the output feature of LLaVA. They also fine-tune the pre-trained diffusion model to enhance the model’s perception and reasoning capability.
@@ 5.1.1 Text to Video Generation with CLIP
Based on the type of generative models, there are two main paradigms: one is based on diffusion models, and the other is based on autoregressive models built with transformer architectures and discrete codebooks, trained with the next-token prediction loss. Diffusion models have become a mainstream paradigm due to their ease of training. Within the diffusion framework, there are pixel-level video diffusion models [12, 16, 17] and latent-level video diffusion models [14, 15, 281, 282, 283, 284, 285, 286]. Pixel-level approaches exhibit better text alignment but require substantial computational resources. On the other hand, latent-level models are more efficient as they reduce redundancy in video data. Different works in video generation have different focuses. Some emphasize photorealistic or high-definition outputs, aiming to improve the quality of generated videos. Others focus on controllable generation, such as image-to-video approaches that enable local control over motion regions, trajectory control for object and camera movement directions, and the use of sketch, depth, and pose to control structure. Some works concentrate on generating longer videos or exploring better network architectures.
@@ 5.1.2 Text to Video Generation with LLMs
Recently, several works have also taken advantage of multimodal LLMs [71, 72] for the task of video generation. For example, VideoPoet [71] leverages a pre-trained auto-regressive transformer model to handle multimodal data for synthesizing videos with temporal consistency and high-motion fidelity. It adapts LLMs training techniques, allowing for task-specific video generation, including tasks such as text-to-video and image-to-video conversion. Another work, MAGVIT-v2 [72], explores the video tokenization technique of MLLMs. It transforms visual inputs into discrete tokens, enhancing the performance of image and video generation tasks. It outperforms diffusion models on benchmarks such as ImageNet and Kinetics, provides video compression on par with advanced codecs, and improves action recognition through effective representation learning.